how much heat is liberated by the combustion of 206 g of hydrogen?
1
Expert's answer
2018-05-08T07:43:27-0400
2H2 + O2 => 2H20
m(H2) = 206g = 0.206kg
q=12*10⁷J/kg
Q - ?
Q = m*q = 0.206 kg* 12*10⁷J/kg = 2.47*10⁷J.